# Estimates for trilinear and quadrilinear character sums

### v3

@@ 6.1. Preliminary transformations
$Q_{1}(d)=\mathop{\sum_{|k|\leqslant 2K/d}\sum_{|n|\leqslant 2N/d}}_{\gcd(ak,bn)=1}$ $\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p\end{subarray}}$
×$\times\sum_{u\sim U}\sum_{w\sim W}\biggl{|}\sum_{v\sim V}\mathrm{e}(v\xi)\chi\left(d\overline{(a\ell_{1}u+bm_{1}w)}(ak\ell_{1}+bm_{1}n)+v\right)$
×$\qquad\qquad\qquad\qquad\times\overline{\chi}\left(d\overline{(a\ell_{2}u+m_{2}w)}(ak\ell_{2}+bm_{2}n)+v\right)\biggr{|}.$
Of course the variables of summation continue to satisfy (6.6). We further impose a restriction that $kw\neq nu$, which introduces an additional error $\mathsf{Err}_{2}(d)$ with
$\mathsf{Err}_{2}(d)\ll KL^{2}M^{2}VWd^{-1}p^{o(1)}.$
Next for $x_{1},x_{2}\in{\mathbb{F}}_{p}^{*}$, $1\leqslant A\ll\mathscr{L}_{2}/d$, $1\leqslant B\ll|a|LU+|b|MW\ll\mathscr{L}_{2}/(dV)$ and $1\leqslant C\leqslant K/d$, we put
$\varrho(x_{1},x_{2})=\varrho(x_{1},x_{2};A,B,C)=\mathop{\sum_{|k|\sim C}\sum_{|n|\leqslant 2N/d}\mathop{\sum\sum}_{\ell_{1},\ell_{2}\leqslant L}\mathop{\sum\sum}_{m_{1},m_{2}\leqslant M}\sum_{u\sim U}\sum_{w\sim W}}_{\begin{subarray}{c}ak\ell_{1}+bm_{1}n\equiv(a\ell_{1}u+bm_{1}w)x_{1}\bmod p\\ ak\ell_{2}+bm_{2}n\equiv(a\ell_{2}u+bm_{2}w)x_{2}\bmod p\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p,\quad kw\neq nu,\quad\gcd(ak,bn)=1\\ |ak\ell_{1}+bm_{1}n|\sim A,\leavevmode\nobreak\ |a\ell_{2}u+bm_{2}w|\sim B\end{subarray}}1,$
where the condition (6.6) continues to apply to the variables of summation. We now have
$Q_{1}(d)\leqslant p^{o(1)}\sup_{\begin{subarray}{c}1\leqslant A\ll\mathscr{L}_{2}/d\\ 1\leqslant B\ll\mathscr{L}_{2}/(dV)\\ 1\leqslant C\leqslant K/d\end{subarray}}\sum_{x_{1},x_{2}\in{\mathbb{F}}_{p}^{*}}\varrho(x_{1},x_{2})\left|\sum_{v\sim V}\mathrm{e}(v\xi)\chi(x_{1}+v)\overline{\chi}(x_{2}+v)\right|+\mathsf{Err}_{2}(d).$
Applying the Hölder inequality with $r\in{\mathbb{N}}$, we find
(6.9) $Q_{1}(d)\leqslant Q_{2}(d)\,p^{o(1)}+\mathsf{Err}_{2}(d)$
@@
(6.10) $Q_{2}(d)=\sup_{\begin{subarray}{c}1\leqslant A\ll\mathscr{L}_{2}/d\\ 1\leqslant B\ll\mathscr{L}_{2}/(dV)\\ 1\leqslant C\leqslant K/d\end{subarray}}\Sigma_{1}^{1-\frac{1}{r}}\left(\Sigma_{2}\cdot\Sigma_{3}\right)^{\frac{1}{2r}}$
@@ 6.2. Bounding $\Sigma_{2}$
$\Sigma_{2}\leqslant(M/C+1)^{2}\left(1+\frac{\mathscr{L}_{2}^{2}}{dpV}\right)^{2}(ABLM)^{1+o(1)}(CW+NU/d)$
for all $1\leqslant A\ll\mathscr{L}_{2}/d$, $1\leqslant B\ll\mathscr{L}_{2}/(dV)$ and $1\leqslant C\leqslant K/d$.
@@
Note that $\Sigma_{2}$ is bounded by the number of tuples
@@
$|k|,|\widetilde{k}|\sim C,$ $\quad\ell_{1},\widetilde{\ell}_{1},\ell_{2},\widetilde{\ell}_{2}\leqslant L,\quad m_{1},\widetilde{m}_{1},m_{2},\widetilde{m}_{2}\leqslant M,\quad|n|,|\widetilde{n}|\leqslant 2N/d,$
$u,\widetilde{u}\sim U,$ $\quad w,\widetilde{w}\sim W,\quad kw\neq nu,\quad\widetilde{k}\widetilde{w}\neq\widetilde{n}\widetilde{u},\quad\gcd(ak,bn)=\gcd(a\widetilde{k},b\widetilde{n})=1,$
$|ak\ell_{1}+bm_{1}n|$ $\sim A,\quad|a\widetilde{k}\widetilde{\ell}_{1}+b\widetilde{m}_{1}\widetilde{n}|\sim A,\quad|a\ell_{2}u+bm_{2}w|\sim B,\quad|a\widetilde{\ell}_{2}\widetilde{u}+b\widetilde{m}_{2}\widetilde{w}|\sim B,$
@@
$p\nmid$ $(a\ell_{1}u+bm_{1}w)(a\ell_{2}u+bm_{2}w)(ak\ell_{1}+bm_{1}n)(ak\ell_{2}+bm_{2}n),$
$p\nmid$ $(a\widetilde{\ell}_{1}\widetilde{u}+b\widetilde{m}_{1}\widetilde{w})(a\widetilde{\ell}_{2}\widetilde{u}+b\widetilde{m}_{2}\widetilde{w})(a\widetilde{k}\widetilde{\ell}_{1}+b\widetilde{m}_{1}\widetilde{n})(a\widetilde{k}\widetilde{\ell}_{2}+b\widetilde{m}_{2}\widetilde{n}),$
@@
(6.13) $(ak\ell_{1}+bm_{1}n)(a\widetilde{\ell}_{1}\widetilde{u}+b\widetilde{m}_{1}\widetilde{w})=(a\widetilde{k}\widetilde{\ell}_{1}+b\widetilde{m}_{1}\widetilde{n})(a\ell_{1}u+bm_{1}w)+z_{1}p,$
@@
(6.14) $(ak\ell_{2}+bm_{2}n)(a\widetilde{\ell}_{2}\widetilde{u}+b\widetilde{m}_{2}\widetilde{w})=(a\widetilde{k}\widetilde{\ell}_{2}+b\widetilde{m}_{2}\widetilde{n})(a\ell_{2}u+bm_{2}w)+z_{2}p,$
with some $0\leqslant|z_{1}|,|z_{2}|\leqslant Z$ where $Z=1+2(|a|KL+|b|MN)(|a|LU+|b|MW)/(dp)$, so by hypothesis we have the inequality
@@
We now fix $m_{1},m_{2},n,k,\ell_{1},\ell_{2},u,w$, and put
$a_{j}=ak\ell_{j}+bm_{j}n,\quad b_{j}=a\ell_{j}u+bm_{j}w$
@@
$|a_{1}|\sim A,\leavevmode\nobreak\ |a_{2}|\leqslant 2\mathscr{L}_{2}/d,\leavevmode\nobreak\ |b_{1}|\leqslant\mathscr{L}_{2}/(dV)\text{ and }|b_{2}|\sim B$
@@
in $x_{1},x_{2},y_{1},y_{2}\in{\mathbb{Z}}$ with $|x_{1}|\leqslant\mathscr{L}_{2}/(dV)$, $|x_{2}|\sim B$, $|y_{1}|\sim A$ and $|y_{2}|\leqslant 2\mathscr{L}_{2}/d$. It is clear that the number $Q$ of such tuples $(x_{1},x_{2},y_{1},y_{2})$ satisfying (6.16) is
(6.17) $Q\ll\left(\frac{A\gcd(a_{1},b_{1})}{a_{1}}+1\right)\left(\frac{B\gcd(a_{2},b_{2})}{b_{2}}+1\right)\ll\gcd(a_{1},b_{1})\gcd(a_{2},b_{2}).$
We now fix one of such tuples $(x_{1},x_{2},y_{1},y_{2})$, satisfying (6.16). Then (6.13) and (6.14) lead to the system of four equations
(6.18) $\begin{cases}a\widetilde{\ell}_{1}\widetilde{u}+b\widetilde{m}_{1}\widetilde{w}=x_{1},\\ a\widetilde{k}\widetilde{\ell}_{1}+b\widetilde{m}_{1}\widetilde{n}=y_{1},\\ a\widetilde{\ell}_{2}\widetilde{u}+b\widetilde{m}_{2}\widetilde{w}=x_{2},\\ a\widetilde{k}\widetilde{\ell}_{2}+b\widetilde{m}_{2}\widetilde{n}=y_{2},\end{cases}$
in $\widetilde{k},\widetilde{\ell}_{1},\widetilde{\ell}_{2},\widetilde{m}_{1},\widetilde{m}_{2},\widetilde{n},\widetilde{u},\widetilde{w}\in{\mathbb{Z}}$. Recalling the conditions ~$\widetilde{\ell}_{1}\widetilde{m}_{2}\not\equiv\widetilde{\ell}_{2}\widetilde{m}_{1}\bmod p$, it suffices to consider the solutions satisfying ~$\widetilde{\ell}_{1}\widetilde{m}_{2}\neq\widetilde{\ell}_{2}\widetilde{m}_{1}$.
We now fix integers $\mathfrak{l}_{1},\mathfrak{l}_{2},\mathfrak{m}_{1},\mathfrak{m}_{2}$ satisfying
@@
(6.20) $a\widetilde{k}\mathfrak{l}_{1}+b\mathfrak{m}_{1}\widetilde{n}=y_{1},\quad a\widetilde{k}\mathfrak{l}_{2}+b\mathfrak{m}_{2}\widetilde{n}=y_{2}$
@@
~$\widetilde{\ell}_{1}=\mathfrak{l}_{1}-s_{1}b\widetilde{n},\quad\widetilde{m}_{1}=\mathfrak{m}_{1}+s_{1}a\widetilde{k},\quad 0\leqslant|s_{1}|\ll M/|\widetilde{k}|\ll M/C,$
@@
~$\widetilde{\ell}_{2}=\mathfrak{l}_{2}-s_{2}b\widetilde{n},\quad\widetilde{m}_{2}=\mathfrak{m}_{2}+s_{2}a\widetilde{k},\quad 0\leqslant|s_{2}|\ll M/|\widetilde{k}|\ll M/C.$
@@
Combining all above arguments, in particular, recalling (6.17), we derive that
$\Sigma_{2}$ $\ll(M/C+1)^{2}Z^{2}\mathop{\sum_{|k|\sim C}\mathop{\sum\sum}_{\ell_{1},\ell_{2}\leqslant L}\mathop{\sum\sum}_{m_{1},m_{2}\leqslant M}\sum_{|n|\leqslant 2N/d}\sum_{u\sim U}\sum_{w\sim W}}_{\begin{subarray}{c}ak\ell_{1}+bm_{1}n=a_{1},\leavevmode\nobreak\ a\ell_{1}u+bm_{1}w=b_{1}\\ ak\ell_{2}+bm_{2}n=a_{2},\leavevmode\nobreak\ a\ell_{2}u+bm_{2}w=b_{2}\\ |a_{1}|\sim A,\leavevmode\nobreak\ |b_{2}|\sim B,\leavevmode\nobreak\ kw\neq nu\end{subarray}}\gcd(a_{1},b_{1})\gcd(a_{2},b_{2}).$
@@
Subsequently, by Lemma 6.1, for all $1\leqslant A\ll\mathscr{L}_{2}/d$, $1\leqslant B\ll|a|LU+|b|MW\ll\mathscr{L}_{2}/(dV)$ and $1\leqslant C\leqslant K/d$, we have the bound
(6.21) $\Sigma_{2}\leqslant(M/C+1)^{2}\left(1+\frac{\mathscr{L}_{2}^{2}}{dpV}\right)^{2}(\mathscr{L}_{2}^{2}LM)^{1+o(1)}(CW+NU/d)d^{-2}V^{-1}.$

### v4

@@ 6.1. Preliminary transformations
$Q_{1}(d)$ $=\mathop{\sum_{|k|\leqslant 2K/d_{1}}\sum_{|n|\leqslant 2N/d_{2}}}_{\gcd(a^{*}k,b^{*}n)=1}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p\end{subarray}}$
×$\qquad\qquad\times\sum_{u\sim U}\sum_{w\sim W}\biggl{|}\sum_{v\sim V}\mathrm{e}(v\xi)\chi\left(\overline{(a\ell_{1}u+bm_{1}w)}(ad_{1}k\ell_{1}+bd_{2}m_{1}n)+v\right)$
×$\qquad\qquad\qquad\qquad\times\overline{\chi}\left(\overline{(a\ell_{2}u+bm_{2}w)}(ad_{1}k\ell_{2}+bd_{2}m_{2}n)+v\right)\biggr{|}.$
Writing $ad_{1}k\ell_{1}+bd_{2}m_{1}n=d(a^{*}k\ell_{1}+b^{*}m_{1}n)$, we now continue as
$Q_{1}(d)$ $=\mathop{\sum_{|k|\leqslant 2K/d_{1}}\sum_{|n|\leqslant 2N/d_{2}}}_{\gcd(a^{*}k,b^{*}n)=1}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p\end{subarray}}$
×$\qquad\times\sum_{u\sim U}\sum_{w\sim W}\biggl{|}\sum_{v\sim V}\mathrm{e}(v\xi)\chi\left(d\overline{(a\ell_{1}u+bm_{1}w)}(a^{*}k\ell_{1}+b^{*}m_{1}n)+v\right)$
×$\qquad\qquad\times\overline{\chi}\left(d\overline{(a\ell_{2}u+bm_{2}w)}(a^{*}k\ell_{2}+b^{*}m_{2}n)+v\right)\biggr{|},$
which can be also interpreted as
$Q_{1}(d)$ $=\mathop{\sum_{|k|\leqslant 2K/d_{1}}\sum_{|n|\leqslant 2N/d_{2}}}_{\gcd(a^{*}k,b^{*}n)=1}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p\end{subarray}}$
×$\qquad\times\sum_{\begin{subarray}{c}u\sim\gcd(a,d)U\\ \gcd(a,d)\mid u\end{subarray}}\sum_{\begin{subarray}{c}w\sim\gcd(b,d)W\\ \gcd(b,d)\mid w\end{subarray}}\biggl{|}\sum_{v\sim V}\mathrm{e}(v\xi)\chi\left(d\overline{(a^{*}\ell_{1}u+b^{*}m_{1}w)}(a^{*}k\ell_{1}+b^{*}m_{1}n)+v\right)$
×$\qquad\qquad\times\overline{\chi}\left(d\overline{(a^{*}\ell_{2}u+b^{*}m_{2}w)}(a^{*}k\ell_{2}+b^{*}m_{2}n)+v\right)\biggr{|}.$
Here we have used the formulas $a^{*}=a/\gcd(a,d)$, $b^{*}=b/\gcd(b,d)$. Of course the variables of summation continue to satisfy (6.6) with necessary changes of variables as in (6.8). We further impose a restriction that $kw\gcd(b,d)\neq nu\gcd(a,d)$, which introduces an additional error $\mathsf{Err}_{2}(d)$ with
$\sum_{d\leqslant D}\mathsf{Err}_{2}(d)\leqslant KL^{2}M^{2}VWp^{o(1)}.$
Next for $x_{1},x_{2}\in{\mathbb{F}}_{p}^{*}$, and for
$1\leqslant A\ll\mathscr{L}_{2}/d,\quad 1\leqslant B\ll|a|LU+|b|MW\ll\mathscr{L}_{2}/V,\quad 1\leqslant C\leqslant K/d_{1},$
we put
(6.10) $\varrho(x_{1},x_{2})=\mathop{\sum_{|k|\sim C}\sum_{|n|\leqslant 2N/d_{2}}\mathop{\sum\sum}_{\ell_{1},\ell_{2}\leqslant L}\mathop{\sum\sum}_{m_{1},m_{2}\leqslant M}\sum_{u\sim U}\sum_{w\sim W}}_{\begin{subarray}{c}a^{*}k\ell_{1}+b^{*}m_{1}n\equiv(a\ell_{1}u+bm_{1}w)x_{1}\bmod p\\ a^{*}k\ell_{2}+b^{*}m_{2}n\equiv(a\ell_{2}u+bm_{2}w)x_{2}\bmod p\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p,\quad kw\gcd(b,d)\neq nu\gcd(a,d),\quad\gcd(a^{*}k,b^{*}n)=1\\ |a^{*}k\ell_{1}+b^{*}m_{1}n|\sim A,\leavevmode\nobreak\ |a\ell_{2}u+bm_{2}w|\sim B\end{subarray}}1,$
where the condition (6.6) continues to apply to the variables of summation with necessary changes of variables as in (6.8). We now have
$Q_{1}(d)\leqslant p^{o(1)}\sup_{\begin{subarray}{c}1\leqslant A\ll\mathscr{L}_{2}/d\\ 1\leqslant B\ll\mathscr{L}_{2}/V\\ 1\leqslant C\leqslant K/d_{1}\end{subarray}}\,\sum_{x_{1},x_{2}\in{\mathbb{F}}_{p}^{*}}\varrho(x_{1},x_{2})\left|\sum_{v\sim V}\mathrm{e}(v\xi)\chi(dx_{1}+v)\overline{\chi}(dx_{2}+v)\right|+\mathsf{Err}_{2}(d).$
Applying the Hölder inequality with a positive integer $r\geqslant 2$, we find
(6.11) $Q_{1}(d)\leqslant Q_{2}(d)\,p^{o(1)}+\mathsf{Err}_{2}(d),$
@@
(6.12) $Q_{2}(d)=\sup_{\begin{subarray}{c}1\leqslant A\ll\mathscr{L}_{2}/d\\ 1\leqslant B\ll\mathscr{L}_{2}/V\\ 1\leqslant C\leqslant K/d_{1}\end{subarray}}\Sigma_{1}^{1-\frac{1}{r}}\left(\Sigma_{2}\cdot\Sigma_{3}\right)^{\frac{1}{2r}}$
@@ 6.2. Bounding $\Sigma_{2}$
$\Sigma_{2}\leqslant(M/C+1)^{2}\left(1+\frac{\mathscr{L}_{2}^{2}}{dpV}\right)^{2}(ABLM)^{1+o(1)}(CW\gcd(b,d)+NU\gcd(a,d)/d_{2})$
for all $1\leqslant A\ll\mathscr{L}_{2}/d$, $1\leqslant B\ll\mathscr{L}_{2}/V$ and $1\leqslant C\leqslant K/d_{1}$.
@@
Let $U_{1}=\gcd(a,d)U$ and $W_{1}=\gcd(b,d)W$. We operate the changes of variables
(6.15) $u\mapsto\frac{u}{\gcd(a,d)},\ \ w\mapsto\frac{w}{\gcd(b,d)}$
in the definition (6.10) of $\rho(x_{1},x_{2})$. Note that $\Sigma_{2}$ is bounded by the number of tuples with length 16
@@
$|k|,|\widetilde{k}|\sim C,$ $\quad\ell_{1},\widetilde{\ell}_{1},\ell_{2},\widetilde{\ell}_{2}\leqslant L,\quad m_{1},\widetilde{m}_{1},m_{2},\widetilde{m}_{2}\leqslant M,\quad|n|,|\widetilde{n}|\leqslant 2N/d_{2},$
$u,\widetilde{u}\sim U_{1},$ $\quad w,\widetilde{w}\sim W_{1},\quad kw\neq nu,\quad\widetilde{k}\widetilde{w}\neq\widetilde{n}\widetilde{u},\quad\gcd(a^{*}k,b^{*}n)=\gcd(a^{*}\widetilde{k},b^{*}\widetilde{n})=1,$
$|a^{*}k\ell_{1}+b^{*}m_{1}n|$ $\sim A,\quad|a^{*}\widetilde{k}\widetilde{\ell}_{1}+b^{*}\widetilde{m}_{1}\widetilde{n}|\sim A,\quad|a^{*}\ell_{2}u+b^{*}m_{2}w|\sim B,\quad|a^{*}\widetilde{\ell}_{2}\widetilde{u}+b^{*}\widetilde{m}_{2}\widetilde{w}|\sim B,$
@@
with the additional non-divisbility conditions
(6.16) $p\nmid(a^{*}\ell_{1}u+b^{*}m_{1}w)(a^{*}\ell_{2}u+b^{*}m_{2}w)(a^{*}k\ell_{1}+b^{*}m_{1}n)(a^{*}k\ell_{2}+b^{*}m_{2}n),$
and
(6.17) $p\nmid(a^{*}\widetilde{\ell}_{1}\widetilde{u}+b^{*}\widetilde{m}_{1}\widetilde{w})(a^{*}\widetilde{\ell}_{2}\widetilde{u}+b^{*}\widetilde{m}_{2}\widetilde{w})(a^{*}\widetilde{k}\widetilde{\ell}_{1}+b^{*}\widetilde{m}_{1}\widetilde{n})(a^{*}\widetilde{k}\widetilde{\ell}_{2}+b^{*}\widetilde{m}_{2}\widetilde{n}),$
@@
(6.18) $(a^{*}k\ell_{1}+b^{*}m_{1}n)(a^{*}\widetilde{\ell}_{1}\widetilde{u}+b^{*}\widetilde{m}_{1}\widetilde{w})=(a^{*}\widetilde{k}\widetilde{\ell}_{1}+b^{*}\widetilde{m}_{1}\widetilde{n})(a^{*}\ell_{1}u+b^{*}m_{1}w)+z_{1}p,$
@@
(6.19) $(a^{*}k\ell_{2}+b^{*}m_{2}n)(a^{*}\widetilde{\ell}_{2}\widetilde{u}+b^{*}\widetilde{m}_{2}\widetilde{w})=(a^{*}\widetilde{k}\widetilde{\ell}_{2}+b^{*}\widetilde{m}_{2}\widetilde{n})(a^{*}\ell_{2}u+b^{*}m_{2}w)+z_{2}p,$
with some $0\leqslant|z_{1}|,|z_{2}|\leqslant Z$ where
$Z=1+4(|a|KL+|b|MN)(|a|LU+|b|MW)/(dp),$
so by hypothesis we have the inequality
@@
Note that the conditions (6.16) and (6.17) are resulted by (6.6) with changes of variables as in (6.8) and (6.15).
We now fix $k,\ell_{1},\ell_{2},m_{1},m_{2},n,u,w$, and put
$a_{j}=a^{*}k\ell_{j}+b^{*}m_{j}n,\quad b_{j}=a^{*}\ell_{j}u+b^{*}m_{j}w$
@@
$|a_{1}|\sim A,\quad|a_{2}|\leqslant 2\mathscr{L}_{2}/d,\quad|b_{1}|\leqslant\mathscr{L}_{2}/V,\quad|b_{2}|\sim B,$
@@
in $x_{1},x_{2},y_{1},y_{2}\in{\mathbb{Z}}$ with $|x_{1}|\leqslant\mathscr{L}_{2}/V$, $|x_{2}|\sim B$, $|y_{1}|\sim A$ and $|y_{2}|\leqslant 2\mathscr{L}_{2}/d$. It is clear that the number of such tuples $(x_{1},x_{2},y_{1},y_{2})$ satisfying (6.21) is bounded, up to an absolute constant, by
(6.22) $\left(\frac{A\gcd(a_{1},b_{1})}{a_{1}}+1\right)\left(\frac{B\gcd(a_{2},b_{2})}{b_{2}}+1\right)\ll\gcd(a_{1},b_{1})\gcd(a_{2},b_{2}).$
We claim that, for given $k,\ell_{1},\ell_{2},m_{1},m_{2},n,u,w,z_{1},z_{2}$ as above, the number ${\mathcal{N}}$ of tuples $(\widetilde{k},\widetilde{\ell}_{1},\widetilde{\ell}_{2},\widetilde{m}_{1},\widetilde{m}_{2},\widetilde{n},\widetilde{u},\widetilde{w})$ satisfying the above-mentioned conditions satisfies
(6.23) ${\mathcal{N}}\ll(M/C+1)^{2}.$
Hence
$\Sigma_{2}$ $\ll(M/C+1)^{2}Z^{2}\mathop{\sum_{|k|\sim C}\mathop{\sum\sum}_{\ell_{1},\ell_{2}\leqslant L}\mathop{\sum\sum}_{m_{1},m_{2}\leqslant M}\sum_{|n|\leqslant 2N/d_{2}}\sum_{u\sim U_{1}}\sum_{w\sim W_{1}}}_{\begin{subarray}{c}a^{*}k\ell_{1}+b^{*}m_{1}n=a_{1},\leavevmode\nobreak\ a^{*}\ell_{1}u+b^{*}m_{1}w=b_{1}\\ a^{*}k\ell_{2}+b^{*}m_{2}n=a_{2},\leavevmode\nobreak\ a^{*}\ell_{2}u+b^{*}m_{2}w=b_{2}\\ |a_{1}|\sim A,\leavevmode\nobreak\ |b_{2}|\sim B,\leavevmode\nobreak\ kw\neq nu\end{subarray}}\gcd(a_{1},b_{1})\gcd(a_{2},b_{2})$
in view of (6.22). Then the desired bound for $\Sigma_{2}$ follows from (6.20) and Lemma 3.3.
It suffices to prove (6.23), and keep in mind that $k,\ell_{1},\ell_{2},m_{1},m_{2},n,u,w,z_{1},z_{2}$ are all fixed. We now fix one of such tuples $(x_{1},x_{2},y_{1},y_{2})$, satisfying (6.21). Then (6.18) and (6.19) lead us to consider the system of four equations
(6.24) $\begin{cases}a^{*}\widetilde{\ell}_{1}\widetilde{u}+b^{*}\widetilde{m}_{1}\widetilde{w}=x_{1},\\ a^{*}\widetilde{k}\widetilde{\ell}_{1}+b^{*}\widetilde{m}_{1}\widetilde{n}=y_{1},\\ a^{*}\widetilde{\ell}_{2}\widetilde{u}+b^{*}\widetilde{m}_{2}\widetilde{w}=x_{2},\\ a^{*}\widetilde{k}\widetilde{\ell}_{2}+b^{*}\widetilde{m}_{2}\widetilde{n}=y_{2},\end{cases}$
in $\widetilde{k},\widetilde{\ell}_{1},\widetilde{\ell}_{2},\widetilde{m}_{1},\widetilde{m}_{2},\widetilde{n},\widetilde{u},\widetilde{w}\in{\mathbb{Z}}$. Recalling the restriction ~$\widetilde{\ell}_{1}\widetilde{m}_{2}\not\equiv\widetilde{\ell}_{2}\widetilde{m}_{1}\bmod p$, it suffices to consider the solutions satisfying ~$\widetilde{\ell}_{1}\widetilde{m}_{2}\neq\widetilde{\ell}_{2}\widetilde{m}_{1}$.
We now fix integers $\mathfrak{l}_{1},\mathfrak{l}_{2},\mathfrak{m}_{1}$ and $\mathfrak{m}_{2}$ satisfying
@@
(6.26) $a^{*}\widetilde{k}\mathfrak{l}_{1}+b^{*}\mathfrak{m}_{1}\widetilde{n}=y_{1},\quad a^{*}\widetilde{k}\mathfrak{l}_{2}+b^{*}\mathfrak{m}_{2}\widetilde{n}=y_{2}$
@@
~$\widetilde{\ell}_{1}=\mathfrak{l}_{1}-s_{1}b^{*}\widetilde{n},\quad\widetilde{m}_{1}=\mathfrak{m}_{1}+s_{1}a^{*}\widetilde{k},\quad 0\leqslant|s_{1}|\ll M/|\widetilde{k}|\ll M/C,$
@@
~$\widetilde{\ell}_{2}=\mathfrak{l}_{2}-s_{2}b^{*}\widetilde{n},\quad\widetilde{m}_{2}=\mathfrak{m}_{2}+s_{2}a^{*}\widetilde{k},\quad 0\leqslant|s_{2}|\ll M/|\widetilde{k}|\ll M/C.$
@@
Subsequently, by Lemma 6.1, for all $1\leqslant A\ll\mathscr{L}_{2}/d$, $1\leqslant B\ll|a|LU+|b|MW\ll\mathscr{L}_{2}/V$ and $1\leqslant C\leqslant K/d_{1}$, we have the bound
(6.27) $\begin{split}\Sigma_{2}\leqslant(M/C+1)^{2}\left(1+\frac{\mathscr{L}_{2}^{2}}{dpV}\right)^{2}&\frac{(\mathscr{L}_{2}^{2}LM)^{1+o(1)}}{dV}\\ &\left(CW\gcd(b,d)+NU\gcd(a,d)/d_{2}\right).\end{split}$